Parth (parthkohli):

Why is my answer wrong? Consider the function on the integers given by \(f(x,y)=x^2y\). How many ordered pairs of integers satisfying \(−16\le x,y\le16\) is \(f(x,y)=f(y,x)\)?

Parth (parthkohli):

\[f(x,y)=f(y,x) \Rightarrow y^2x = x^2 y \]I observed that \(x\) and \(y\) cannot be negative here. So I am left with numbers ranging from \(0\) to \(16\). That gives me \(17\).

Parth (parthkohli):

Are there any other ordered pairs where \(x \ne y\)?

Parth (parthkohli):

ZOMG, wait.

Parth (parthkohli):

\((-4)^2 \times -4 = (-4)^2 \times -4\)

Parth (parthkohli):

Facepalm. They can be negative.

OpenStudy (anonymous):

PING and the light turns on!

OpenStudy (anonymous):

:D

Parth (parthkohli):

lol yeah. That fetches me \(17 + 16 = 33\)

Parth (parthkohli):

Is 33 correct?

OpenStudy (anonymous):

no

Parth (parthkohli):

Aww.

OpenStudy (anonymous):

you have 33+16+16=?

Parth (parthkohli):

Why is that so?

OpenStudy (anonymous):

that many ordered pairs (0,?) and (?,0)

Parth (parthkohli):

Oh, I'm stupid. Thanks man!

OpenStudy (anonymous):

so, it'd be 33+32+32 ordered pairs

OpenStudy (anonymous):

how many ways can you have 1) x=0 2) y=0 3) x=y

Parth (parthkohli):

(1,1),(2,2)...(16,16) is 16. (-1,-1)...(-16,-16) is 16. (0,?) is 33. (?,0) is 33. 32 + 33 + 33 = 98 We overcounted a (0,0), so 32 + 32 + 33 = 97. Right?

OpenStudy (anonymous):

Is this from brilliant.org?

Parth (parthkohli):

@oldrin.bataku Yeah, practicing stuff :-)

OpenStudy (anonymous):

this way you counted (0,0) twice

Parth (parthkohli):

Am I not allowed to ask questions for practice?

Parth (parthkohli):

@electrokid I subtracted it in the end.

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

I'm not sure if you're allowed to ask for help on brilliant.org questions, at least publicly... :-p

OpenStudy (perl):

yes you are allowed

OpenStudy (perl):

and i am god

Parth (parthkohli):

@oldrin.bataku But I think I'm just practicing techniques.

OpenStudy (perl):

brilliant! 97 is the correct answer

OpenStudy (anonymous):

o'course it'd be lol

OpenStudy (perl):

u wbhi dheu dhee hee:)

OpenStudy (anonymous):

@perl what?

Parth (parthkohli):

@perl ?

OpenStudy (perl):

i said, i enjoyed the problem

OpenStudy (perl):

now i must go , there are babies to punish